# Thomassen’s proof and Filippov’s proof of the Weak Jordan Theorem

### 05 Apr 2026 v1

@@ 2 Thomassen’s proof of the WJT
Step 1. The JT holds for PL circuits. Suppose that $f\colon I\to\mathbb{R}^{2}$ is a PL circuit. We first prove that the open set $D=\mathbb{R}^{2}\setminus f[I]$ (Proposition 1.6) is disconnected. We define a continuous map $g\colon D\to\mathbb{R}$ such that $g[D]=\{0,1\}$. By Proposition 1.11, the set $D$ is disconnected because $\{0,1\}$ is a disconnected subset of $\mathbb{R}$.
@@
where << is the down-up order and every $z_{i}$ is either a union of consecutive (vertical) segments of $f[I]$ or a point in $f[I]$. The $z_{i}$ are mutually disjoint. We call $z_{i}$simple intersection if locally near $z_{i}$, the set $f[I]$ lies on both sides of $\ell(a)$. Else, if locally $f[I]$ lies on only one side of $\ell(a)$, we call $z_{i}$double intersection. We define $g(a)\in\{0,1\}$ as
@@
We show that $g$ is continuous. Let $a\in D$. We take a ball $B=B(a,r)\subset D$ with radius $r>0$ so small that no half-line $\ell(a^{\prime})$ with $a^{\prime}\in B$ contains a corner of $f[I]$ not present in $P(a)$. Now if we move $\ell(a)$ to $\ell(a^{\prime})$ ($a^{\prime}\in B$), simple intersections in $P(a)$ are preserved, and every double intersection either does not change (when $\ell(a^{\prime})\subset\ell(a)$ or $\ell(a)\subset\ell(a^{\prime})$) or disappears or decays into two simple (point) intersections. Importantly, no new intersections besides those born from $P(a)$ appear in $P(a^{\prime})$. It follows that the number of simple intersections increases by an even number in $\mathbb{N}_{0}$, and therefore $g(a)=g(a^{\prime})$. The function $g$ is continuous at $a$ and hence on $D$.
By Proposition 1.4, $f[I]$ is a compact set. By Proposition 1.5, $f[I]$ is closed and bounded. It follows that $f[I]$ has a highest point, which in fact is a corner. Considering the values of $g$ near this highest corner, we see that $g[D]=\{0,1\}$.
@@
$D=\mathbb{R}^{2}\setminus f[I]=L(f)\cup R(f)\,$,
where $L(f)$ and $R(f)$ are nonempty open connected sets that are disjoint. To define these sets, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of the domain $I=[a,b]$ of $f$ and list the corners $c_{i}=f(a_{i})$ and segments $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ for $i=1,2,\dots,k$ in cyclic (indices are taken modulo $k$) order
@@
We fix one of the two consistent orientations of the segments $s_{i}$ and define $L(f)$ as the set of points $c\in D$ that can be joined to $f[I]$ in $D$ from the left. It means that there is a PL map $f_{c}\colon I=[a,b]\to\mathbb{R}^{2}$ such that $f_{c}(a)=c$, $f_{c}(b)\in f[I]$, $f_{c}^{0}\subset D$, and the last segment of $f_{c}[I]$ approaches $f[I]$ from the left side with respect to the fixed orientation. Note that $f_{c}^{0}\subset L(f)$. We similarly define $R(f)$, by joining $c$ to $f[I]$ in $D$ from the right. It is easy to see that the sets $L(f)$ and $R(f)$ are nonempty and open, and that $L(f)\cup R(f)=D$. If they are connected, it follows that they are disjoint: if $L(f)\cap R(f)\neq\emptyset$, then $L(f)\cup R(f)=D$ is connected by Proposition 1.13, which contradicts the previous result.
So it remains to prove that $L(f)$ and $R(f)$ are connected sets; the remaining properties stated in the JT, namely that one of them is bounded, the other is unbounded, and that $f[I]$ is their common boundary, follow easily.
We prove that $L(f)$ is connected; the argument for $R(f)$ is similar. We fix a point $c\in L(f)$ and for every point $d\in L(f)$ we define a connected set $C_{d}$ such that $c,d\in C_{d}\subset L(f)$. Then
@@
We define sets $C_{d}$. For $\delta>0$ and $i=1,2,\dots,k$, we denote by $\ell_{i}$ ($\ni c_{i}$) the axis of the two angles determined by $s_{i}$ and $s_{i+1}$. We denote by $c_{i}(\delta)\in\ell_{i}$ the point on $\ell_{i}$ lying in a distance of $\delta$ from $c_{i}$ on the left of $f[I]$. Let $s_{i}(\delta)$ be the straight segment joining $c_{i-1}(\delta)$ and $c_{i}(\delta)$. Clearly,
$s_{i}(\delta)\cap s_{j}=\emptyset,\ j=i-1,\,i,\,i+1\,$,
and $d_{2}(c,s_{i})<\delta$ for every $c\in s_{i}(\delta)$. It follows that there is a $\delta_{0}$ such that $s_{i}(\delta)\subset L(f)$ for every $\delta\leq\delta_{0}$. Let
@@
Then $C(\delta)\subset L(f)$ for every sufficiently small $\delta$. Every set $C(\delta)$ is connected by Proposition 1.11 (which implies that every segment $s_{i}(\delta)$ is connected) and Proposition 1.13 (the intersection graph spans a $k$-cycle). Let $c,d\in L(f)$ be the above mentioned points, and let $f_{c}$ and $f_{d}$ be the PL maps that join them to $f[I]$ in $D$ from the left. We take small enough $\delta$ such that $C(\delta)\subset L(f)$ and that
@@
is the desired set. It is connected by Propositions 1.11 and 1.13.
Step 2. PL $K_{3,3}$-configurations do not exist. We assume that $U$, $V$ ($\subset\mathbb{R}^{2}$) and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a PL $K_{3,3}$-configuration and obtain a contradiction. We may assume that $U=\{1,3,5\}$ and $V=\{2,4,6\}$. We consider the PL circuit $C\colon I\to\mathbb{R}^{2}$ given by the $6$-cycle
@@
We fix one of the two orientations of $C[I]$ and consider the partition $\mathbb{R}^{2}\setminus C[I]=D=L(C)\cup R(C)$ of the complement defined in the previous step. For the three remaining arcs
@@
each set $e_{i}^{0}$ lies completely either in $L(C)$ or in $R(C)$ ($e_{i}^{0}$ is connected by Proposition 1.11). Thus, two of the $e_{i}^{0}$, for example, $e_{2}^{0}$ and $e_{3}^{0}$, lie in one set, for example, $R(C)$. We consider the PL circuit $C^{\prime}\colon I\to\mathbb{R}^{2}$ given by the $4$-cycle
@@
We orient $C^{\prime}[I]$ consistently with the chosen orientation of $C[I]$ (the first three arcs of $C^{\prime}$ already have this orientation). Let $g=e_{3}+f_{6,5}$. We consider the partition
$\mathbb{R}^{2}\setminus C^{\prime}[I]=D^{\prime}=L(C^{\prime})\cup R(C^{\prime})\,.$
Near the vertex $5$, the set $g^{0}$ ($\subset D^{\prime}$) intersects $L(C^{\prime})$. Near the vertex $3$, it intersects $R(C^{\prime})$. Hence $L(C^{\prime})$ and $R(C^{\prime})$ cut $g^{0}$, which is a contradiction because $g^{0}$ is connected by Proposition 1.11. In each of the five other cases (when two $e_{i}^{0}$ lie in $L(C^{\prime})$ or in $R(C^{\prime})$) we obtain a similar contradiction.
Step 3. $K_{3,3}$-configurations do not exist. We assume that $U$, $V$ and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a $K_{3,3}$-configuration. We transform it in a PL $K_{3,3}$-configuration, which by the previous step is a contradiction. Let $U=\{u_{1},u_{2},u_{3}\}$ and $V=\{u_{4},u_{5},u_{6}\}$. To simplify notation, for $1\leq i\leq 3$ and $4\leq j\leq 6$ we write $f_{i,j}$ instead of $f_{u_{i},u_{j}}$. For $i=1,2,\dots,6$, let $B_{i}=\overline{B}(u_{i},r)$ be mutually disjoint closed balls with the radius $r>0$ so small that $f_{i,j}[I]\cap B_{l}=\emptyset$ whenever $l\neq i,j$ (Proposition 1.8). Let $D=\mathbb{R}^{2}\setminus\bigcup_{i=1}^{6}B_{i}$. Using the arcs $f_{i,j}$ and Proposition 1.7, we obtain nine arcs $g_{i,j}\colon I=[a,b]\to\mathbb{R}^{2}$ such that $g_{i,j}$ joins the circle $\partial B_{i}$ to the circle $\partial B_{j}$, that $g_{i,j}^{0}\subset D$ and that the nine sets $g_{i,j}[I]$ are mutually disjoint. Using Propositions 1.8 and 1.10, we obtain $\delta>0$ and nine partitions of $[a,b]$,
$P_{i,\,j}=(a=a_{0,\,i,\,j}<a_{1,\,i,\,j}<\dots<a_{k(i,\,j),\,i,\,j}=b)\,$,
@@
$g(a_{l-1,\,i,\,j})\in B(g(a_{l,\,i,\,j}),\,\delta)\,\text{ for }\,l=1,\,2,\,\dots,\,k(i,\,j)\,$,
that
$B(g(a_{l,\,i,\,j}),\,\delta)\cap B(g(a_{l^{\prime},\,i^{\prime},\,j^{\prime}}),\,\delta)=\emptyset$
for every $l=0,1,\dots,k(i,j)$, $l^{\prime}=0,1,\dots,k(i^{\prime},j^{\prime})$ and every pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$, and that $u_{r}\not\in B(g(a_{l,i,j}),\delta)$ for every $r=1,2,\dots,6$ and $l=0,1,\dots,k(i,j)$. Let
$U_{i,\,j}=\bigcup_{l=0}^{k(i,\,j)}B(g(a_{l,\,i,\,j}),\,\delta)\,.$}
Each of these nine open sets is connected by Propositions 1.13 and 1.14 (the intersection graph spans a path), and they are mutually disjoint. Using Proposition 1.15, we obtain nine PL arcs $h_{i,j}\colon[a,b]\to U_{i,j}$ such that $h_{i,j}$ joins $\alpha_{i,j}=h_{i,j}(a)=g_{i,j}(a)\in\partial B_{i}$ to $\beta_{i,j}=h_{i,j}(b)=g_{i,j}(b)\in\partial B_{j}$ in $U_{i,j}$. We consider the straight segments $u_{i}\,\alpha_{i,j}$ and $\beta_{i,j}\,u_{j}$ and easily modify the PL arcs $h_{i,j}$ to nine PL arcs $p_{i,j}$ such that $p_{i,j}$ joins $u_{i}$ to $u_{j}$ and
@@
for every pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$. Then $U$, $V$ and $\{p_{i,j}\colon\;1\leq i\leq 3,\,4\leq j\leq 6\}$ is a PL $K_{3,3}$-configuration.
@@
So let $f\colon I\to\mathbb{R}^{2}$ be a circuit with the property that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is connected. Let $a$ and $b$ be points in $f[I]$ with the maximum and minimum $y$-coordinate, respectively. These points exist by Propositions 1.4 and 1.5. Clearly, $a_{y}>b_{y}$ (Proposition 1.16). By Proposition 1.17 the points $a$ and $b$ divide $f[I]$ in two arcs $f_{0},f_{1}\colon I\to\mathbb{R}^{2}$ such that $f_{0}^{0}\cap f_{1}^{0}=\emptyset$, $f_{0}[I]\cup f_{1}[I]=f[I]$, and $\mathrm{ep}(f_{0})=\mathrm{ep}(f_{1})=\{a,b\}$. Let $f_{2}\colon I\to\mathbb{R}^{2}$ be a PL arc joining $a$ to $b$ such that $f_{2}^{0}\subset D$. For example, $f_{2}$ starts with a vertical segment going up from $a$, then goes horizontally sufficiently far to the right, then goes vertically down to a level below $b_{y}$, then goes horizontally left just below $b$, and closes by going vertically up to $b$. Such PL arc $f_{2}$ exists by Propositions 1.4 and 1.5. Let $\ell$ be any horizontal line with $y$-coordinate in $(b_{y},a_{y})$. By Propositions 1.18 and 1.19, the line $\ell$ contains a straight segment $T$ such that one endpoint $c$ is in $f_{0}^{0}$, the other endpoint $d$ is in $f_{1}^{0}$ and $T^{0}\subset D$. By the connectedness of $D$ and by Propositions 1.6, 1.7 and 1.15, there exists an arc $f_{e,g}$ such that it joins a point $e\in T^{0}$ to a point $g\in f_{2}^{0}$, $f_{e,g}[I]\subset D$ and $f_{e,g}^{0}\cap(T\cup f_{2}[I])=\emptyset$. We set $U=\{c,d,g\}$, $V=\{a,b,e\}$ and define the connecting arcs. The arc $f_{e,g}$ is already defined. The arcs $f_{e,c}$ and $f_{e,d}$ are the corresponding subsegments of $T$. The arcs $f_{a,c}$ and $f_{a,d}$ are the corresponding subarcs of $f_{0}$ and $f_{1}$, respectively. The arc $f_{a,g}$ is the corresponding subarc of $f_{2}$. The arcs $f_{b,c}$, $f_{b,d}$, and $f_{b,g}$ are defined similarly to those starting at $a$. It is easy to check that the nine defined arcs have mutually disjoint interiors. We get the $K_{3,3}$-configuration $U$, $V$ and $\{f_{u,v}\colon\;u\in U,v\in V\}$.

### 19 Apr 2026 v3

@@ 2 Thomassen’s proof of the WJT
Claim 1. The JT holds for PL circuits. Suppose that $f\colon I\to\mathbb{R}^{2}$ is a PL circuit. We first prove that the open set $D=\mathbb{R}^{2}\setminus f[I]$ (Proposition 1.8) is disconnected: we define a continuous map $g\colon D\to\mathbb{R}$ such that $g[D]=\{0,1\}$. By Proposition 1.13, the set $D$ is disconnected because $\{0,1\}$ is a disconnected subset of $\mathbb{R}$.
@@
where << is the vertical order and every $z_{i}$ is either a union of consecutive (vertical) segments of $f[I]$ or a point in $f[I]$. The $z_{i}$ are mutually disjoint. We call $z_{i}$simple intersection if locally near $z_{i}$, the set $f[I]$ lies on both sides of $\ell(a)$. Else, if locally $f[I]$ lies on only one side of $\ell(a)$, we call $z_{i}$double intersection. We define $g(a)\in\{0,1\}$ as
@@
We show that $g$ is continuous. Let $a\in D$. We take a ball $B=B(a,r)\subset D$ with radius $r>0$ so small that no half-line $\ell(a^{\prime})$ with $a^{\prime}\in B$ contains a corner of $f[I]$ not present in $P(a)$. If we move $a$ to $a^{\prime}\in B$, simple intersections in $P(a)$ are preserved, and every double intersection either does not change (when $\ell(a^{\prime})\subset\ell(a)$ or $\ell(a)\subset\ell(a^{\prime})$) or disappears or decays into two simple (point) intersections. Importantly, no new intersections besides those born from $P(a)$ are introduced. It follows that the number of simple intersections increases by an even number in $\mathbb{N}_{0}=\{0,1,\dots\}$, and therefore $g(a)=g(a^{\prime})$. The function $g$ is continuous at $a$ and on $D$.
Since $f[I]$ is a union of finitely many straight segments, it follows that $f[I]$ has a highest point, which is a corner. Considering the values of $g$ near this highest corner, we see that $g[D]=\{0,1\}$. This concludes the proof that $D$ is disconnected.
@@
$D=\mathbb{R}^{2}\setminus f[I]=L(f)\cup R(f)$
for two nonempty open connected and disjoint sets $L(f)$ and $R(f)$. To define them, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of the domain $I=[a,b]$ of $f$ and list the corners $c_{i}=f(a_{i})$ and segments $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ for $i=1,2,\dots,k$ in cyclic (indices are taken modulo $k$) order
@@
We fix one of the two orientations of $f[I]$ and define $L(f)$ as the set of points $c\in D$ that can be joined to $f[I]$ in $D$ from the left. It means that there is a PL map $f_{c}\colon I=[a,b]\to\mathbb{R}^{2}$ such that $f_{c}(a)=c$, $f_{c}(b)\in f[I]$, $f_{c}^{0}\subset D$, and the last segment of $f_{c}[I]$ approaches $f[I]$ from the left side with respect to the fixed orientation. Note that $f_{c}^{0}\subset L(f)$. We similarly define $R(f)$, by joining $c\in D$ to $f[I]$ in $D$ from the right. It is easy to see that the sets $L(f)$ and $R(f)$ are nonempty and open (Lemma 1.17 and its proof), and that $L(f)\cup R(f)=D$. If they are connected, it follows that they are disjoint: if $L(f)\cap R(f)\neq\emptyset$, then $L(f)\cup R(f)=D$ is connected by Proposition 1.15, which contradicts the previous result.
So it remains to prove that $L(f)$ and $R(f)$ are connected sets. The remaining properties stated in the JT, namely that one of the sets is bounded, the other is unbounded, and that $f[I]$ is the common boundary, follow easily.
We prove that $L(f)$ is connected (the argument for $R(f)$ is similar). We fix a point $c\in L(f)$ and for every point $d\in L(f)$ we define a connected set $C_{d}$ such that $c,d\in C_{d}\subset L(f)$. Then
@@
We define sets $C_{d}$. For $i=1,2,\dots,k$, we denote by $\ell_{i}$ ($\ni c_{i}$) the axis of the two angles determined by $s_{i}$ and $s_{i+1}$. For $\delta>0$ let $c_{i}(\delta)\in\ell_{i}$ be the point on $\ell_{i}$ lying in a distance of $\delta$ from $c_{i}$ on the left of $f[I]$. Let $s_{i}(\delta)$ be the straight segment joining $c_{i-1}(\delta)$ and $c_{i}(\delta)$. If $\delta$ is sufficiently small, then
$s_{i}(\delta)\cap s_{j}=\emptyset\,\text{ for }\,j=i-1,\,i,\,i+1\,.$
Also, $d_{2}(c,s_{i})\leq\delta$ for every $c\in s_{i}(\delta)$. It follows (Propositions 1.6 and 1.10) that there is a $\delta_{0}>0$ such that $s_{i}(\delta)\subset L(f)$ for every $\delta\leq\delta_{0}$. Let
@@
Then $C(\delta)\subset L(f)$ for every sufficiently small $\delta$. Every set $C(\delta)$ is connected by Proposition 1.13 (which implies that every segment $s_{i}(\delta)$ is connected) and Proposition 1.15 (the intersection graph spans a $k$-cycle). Let $c,d\in L(f)$ be the above mentioned points, and let $f_{c}$ and $f_{d}$ be respective PL maps that join them to $f[I]$ in $D$ from the left. We take small enough $\delta$ such that $C(\delta)\subset L(f)$ and that
@@
is the desired set. It is connected by Propositions 1.13 and 1.15.
Claim 2. PL $K_{3,3}$-configurations do not exist. We assume that $U$, $V$ ($\subset\mathbb{R}^{2}$) and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a PL $K_{3,3}$-configuration and obtain a contradiction. Let $U=\{u_{1},u_{3},u_{5}\}$ and $V=\{u_{2},u_{4},u_{6}\}$. We write $f_{i,j}$ instead of $f_{u_{i},u_{j}}$. We consider the PL circuit $C\colon I\to\mathbb{R}^{2}$ given by the $6$-cycle
@@
We fix one of the two orientations of $C[I]$ and consider the partition $D=\mathbb{R}^{2}\setminus C[I]=L(C)\cup R(C)$ of $D$ defined in Step 1. Let
@@
be the three remaining arcs of the $K_{3,3}$-configuration. Each set $e_{i}^{0}$ lies completely either in $L(C)$ or in $R(C)$ ($e_{i}^{0}$ is connected by Proposition 1.13). Two of them, for example, $e_{2}^{0}$ and $e_{3}^{0}$, lie in one set, for example, $R(C)$. We consider the PL circuit $C^{\prime}\colon I\to\mathbb{R}^{2}$ given by the $4$-cycle
@@
We orient $C^{\prime}[I]$ in accordance with the chosen orientation of $C[I]$ (the first three arcs of $C^{\prime}$ already have this orientation). Let $g=e_{3}+f_{6,5}$. We consider the partition
$D^{\prime}=\mathbb{R}^{2}\setminus C^{\prime}[I]=L(C^{\prime})\cup R(C^{\prime})\,.$
Near the corner $u_{5}$, the set $g^{0}$ ($\subset D^{\prime}$) intersects $L(C^{\prime})$. Near the corner $u_{3}$, $g^{0}$ intersects $R(C^{\prime})$. Hence $L(C^{\prime})$ and $R(C^{\prime})$ cut $g^{0}$, which is a contradiction because $g^{0}$ is connected by Proposition 1.13. In each of the five other cases (when two $e_{i}^{0}$ lie in $L(C^{\prime})$ or in $R(C^{\prime})$) we obtain a similar contradiction.
Claim 3. $K_{3,3}$-configurations do not exist. We assume that $U$, $V$, and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a $K_{3,3}$-configuration. We transform it in a PL $K_{3,3}$-configuration, which by Claim 2 is a contradiction. Let $U=\{u_{1},u_{2},u_{3}\}$ and $V=\{u_{4},u_{5},u_{6}\}$. We again write $f_{i,j}$ instead of $f_{u_{i},u_{j}}$. For $i=1,2,\dots,6$, let
$B_{i}=\overline{B}(u_{i},\,r)=\{x\in\mathbb{R}^{2}\colon\;d_{2}(x,\,u_{i})\leq r\}$
be mutually disjoint closed balls with the radius $r>0$ so small that $f_{i,j}[I]\cap B_{l}=\emptyset$ whenever $l\neq i,j$ (Propositions 1.6 and 1.10). Let $D=\mathbb{R}^{2}\setminus\bigcup_{i=1}^{6}B_{i}$. Using the arcs $f_{i,j}$ and Proposition 1.9, we obtain nine subarcs $g_{i,j}\colon I=[a,b]\to\mathbb{R}^{2}$ of $f_{i,j}$ such that $g_{i,j}$ joins a point $\alpha_{i,j}\in\partial B_{i}$ to a point $\beta_{i,j}\in\partial B_{j}$, that $g_{i,j}^{0}\subset D$ and that the nine sets $g_{i,j}[I]$ are mutually disjoint. Using Propositions 1.6, 1.10, and 1.12, we obtain $\delta>0$ and nine partitions of $[a,b]$,
$a=a_{0,\,i,\,j}<a_{1,\,i,\,j}<\dots<a_{k(i,\,j),\,i,\,j}=b\ \ (k(i,\,j)\in\mathbb{N}=\{1,2,\dots\})\,$,
@@
$g_{i,\,j}(a_{l-1,\,i,\,j})\in B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\,\text{ for }\,l=1,\,2,\,\dots,\,k(i,\,j)\,$,
and that
$B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\cap B(g_{i^{\prime},\,j^{\prime}}(a_{l^{\prime},\,i^{\prime},\,j^{\prime}}),\,\delta)=\emptyset$
for every $l=0,1,\dots,k(i,j)$, $l^{\prime}=0,1,\dots,k(i^{\prime},j^{\prime})$ and pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$. Let
$U_{i,\,j}=\bigcup_{l=0}^{k(i,\,j)}B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\,.$}
Each of these nine open sets is connected by Propositions 1.15 and 1.16 (the intersection graph spans a path), and they are mutually disjoint. Using Proposition 1.18, we obtain nine PL arcs $h_{i,j}\colon[a,b]\to U_{i,j}$ such that $h_{i,j}$ joins $\alpha_{i,j}=h_{i,j}(a)=g_{i,j}(a)$ ($\in\partial B_{i}$) to $\beta_{i,j}=h_{i,j}(b)=g_{i,j}(b)$ ($\in\partial B_{j}$) in $U_{i,j}$. We consider the first intersections of the straight segments $u_{i}\,\alpha_{i,j}$ and $u_{j}\,\beta_{i,j}$ with $h_{i,j}[I]$ (Proposition 1.9) and obtain nine PL arcs $p_{i,j}$ such that $p_{i,j}$ joins $u_{i}$ to $u_{j}$ and
@@
for every pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$. Then $U$, $V$, and $\{p_{i,j}\colon\;1\leq i\leq 3,\,4\leq j\leq 6\}$ is a PL $K_{3,3}$-configuration.
@@
So let $f\colon I\to\mathbb{R}^{2}$ be a circuit with the property that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is connected. Let $a$ and $b$ be points in $f[I]$ with the maximum and minimum $y$-coordinate, respectively (Propositions 1.6 and 1.19). Clearly, $a_{y}>b_{y}$ (Proposition 1.20). By Proposition 1.21 the points $a$ and $b$ divide $f[I]$ in two arcs $f_{0},f_{1}\colon I\to\mathbb{R}^{2}$ such that $f_{0}^{0}\cap f_{1}^{0}=\emptyset$, $f_{0}[I]\cup f_{1}[I]=f[I]$, and $\mathrm{ep}(f_{0})=\mathrm{ep}(f_{1})=\{a,b\}$. Let $f_{2}\colon I\to\mathbb{R}^{2}$ be a PL arc joining $a$ to $b$ such that $f_{2}^{0}\subset D$. For example, $f_{2}$ starts with a vertical segment going up from $a$, then goes horizontally sufficiently far to the right, goes vertically down to a level below $b_{y}$, continues horizontally left just below $b$, and closes by going vertically up to $b$. Such a PL arc $f_{2}$ exists by Propositions 1.6 and 1.7. Let $\ell$ be any horizontal line with $y$-coordinate in $(b_{y},a_{y})$. By Propositions 1.22 and 1.23, the line $\ell$ contains a straight segment $T$ such that one endpoint $c$ is in $f_{0}^{0}$, the other endpoint $d$ is in $f_{1}^{0}$ and $T^{0}\subset D$. By the connectedness of $D$ and by Propositions 1.8, 1.9 and 1.18, there exists an arc $f_{e,g}$ joining a point $e\in T^{0}$ to a point $g\in f_{2}^{0}$ in $D$ and such that $f_{e,g}^{0}\cap(T\cup f_{2}[I])=\emptyset$. We set $U=\{a,b,e\}$, $V=\{c,d,g\}$ and define the connecting arcs. The arc $f_{e,g}$ is already defined. The arcs $f_{e,c}$ and $f_{e,d}$ are the corresponding subsegments of $T$. The arcs $f_{a,c}$ and $f_{a,d}$ are the corresponding subarcs of $f_{0}$ and $f_{1}$, respectively. The arc $f_{a,g}$ is the corresponding subarc of $f_{2}$. The arcs $f_{b,c}$, $f_{b,d}$, and $f_{b,g}$ are defined similarly to those starting at $a$. It is easy to check that the nine arcs defined have mutually disjoint interiors and that $U$, $V$, and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a $K_{3,3}$-configuration.